OpenStudy (anonymous):

Please use substitution to solve this system of equations. 2x-3y=-1 y=x-1 A. (-5,1) B.(4,3) C.(-2,-3) D.(5,-3)

OpenStudy (anonymous):

Can't you do this?

OpenStudy (anonymous):

No... I always get decimals..

OpenStudy (anonymous):

show me your steps.

OpenStudy (anonymous):

alright. Let me type them

OpenStudy (anonymous):

2x-3y=-1 y=x-1 2(-5) - 3(1) = -1 -13 2(4) - 3(3) = -1 -1 2(-2)-3(-3)=-1 5 2(5)-3(-3) =-1 19 (whew. i used a calc so i wouldn't mess up. idk how to do it..)

OpenStudy (anonymous):

No this problem you can solve it,so you don't need to waste your time substituting in it. - * - = + - * + = -